OpenStudy (anonymous):

write polynomial functions with these features. a) a linear function whose graph has x-intercepts 4. Please help!!

OpenStudy (anonymous):

x-intercepts? plural?

OpenStudy (anonymous):

Only 1 x-intercept?

OpenStudy (anonymous):

sorry its singular x intercept of 4

OpenStudy (mertsj):

x=4

OpenStudy (anonymous):

And what degree polynomial are you instructed to write?

OpenStudy (mertsj):

it says linear.

OpenStudy (anonymous):

a linear function i guess

OpenStudy (anonymous):

But x=4, therefore (x-4) is one linear factor

OpenStudy (mertsj):

That would be degree 1

OpenStudy (anonymous):

So according to the given information, you can make up another one

OpenStudy (anonymous):

You cannot have a linear function with more than one x-intercept.

OpenStudy (anonymous):

linear is y=mx+b right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would it be y=4x+0?

OpenStudy (anonymous):

So x-4=0

OpenStudy (anonymous):

Yes, but what would be the function? y = ?

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Or f(x)=x-4

OpenStudy (anonymous):

oh so where does the b go?

OpenStudy (anonymous):

@pretyroze Do you understand that?

OpenStudy (anonymous):

b is the constant 4

OpenStudy (anonymous):

i thought b was x?

OpenStudy (anonymous):

@pretyroze b = 0, so y = 4x + 0 which simplifies to y = 4x.

OpenStudy (anonymous):

No x is x, and m is the number before x, so in 2x-3 for example, 2 is m, x is x, and -3 is b

OpenStudy (anonymous):

@pretyroze b is a contant, it represents the y-intercept of the straight line.

OpenStudy (anonymous):

that makes a lot of sense

OpenStudy (anonymous):

what if the question asks for a quadratic function whose graph has only one x intercept of 4? how would i set that up?

OpenStudy (anonymous):

@pretyroze ...sorry, the equation is y = x -4...that has an x-intercept of x = 4. so in that equation, -4 is the y-intercept.

OpenStudy (anonymous):

why do you subtract 4?

OpenStudy (anonymous):



OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

Look at the grap of y = x -4. You can see that x = 4 is an x-intercept.

OpenStudy (anonymous):

yes and the y is a negative 4

OpenStudy (anonymous):

yes, y-intercept of - 4.

OpenStudy (anonymous):

what if the question asks for a quadratic function whose graph has only one x intercept of 4? how would i set that up?

OpenStudy (anonymous):

You would have to make up the second intercept because a quadratic always has 2

OpenStudy (anonymous):

(x-4)(x+2) and then factor

OpenStudy (anonymous):

so would it be the conjugate of 4 which is -4?

OpenStudy (anonymous):

Or whatever number in substitute of 2

OpenStudy (anonymous):

where did the 2 come from

OpenStudy (anonymous):

Made it up

OpenStudy (anonymous):

You can't have a definite quadratic with only one x-intercept

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

so i can plug in any number?

OpenStudy (anonymous):

So if they said the zero (x-intercept) is 3, than you would have to make up the other

OpenStudy (anonymous):

But if they said it is 1+3i, (a complex zero) than you would have to multiply it by it's conjugate, (1-3i)

OpenStudy (anonymous):

But yes, you can plug in any number. As long as only 1 x-intercept is given

OpenStudy (anonymous):

ohhh so conjugates are only used with complex numbers?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

Ex. 1+5i. 1-5i 3+2i, 3-2i

OpenStudy (anonymous):

And they are factored like this (x-(1+3i))(x-(1-3i))

OpenStudy (anonymous):

Conjugates are also with irrational numbers, 2 + sqrt(3) and 2 - sqrt(3).

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Not just for comples numbers.

OpenStudy (anonymous):

Forgot about that ^

OpenStudy (anonymous):

Also to remember, every degree up the polynomial goes, the number of zeros does as well. Linear: 1 Quadratic: 2 3rd degree: 3 4th degree: you get the point

OpenStudy (anonymous):

what if it asks us to do the same thing but for a cubic functon whose graph has only one x intercept of 4. I know that a cubic function is y=a(x-r1)(x-r2)(x-r3)

OpenStudy (anonymous):

Then you have to make up 2 zeros

OpenStudy (anonymous):

You probably won't get a question like that though

OpenStudy (anonymous):

A cubic function means 3rd degree polynomial, therefore it has 3 zeros (x-intercepts)

OpenStudy (anonymous):

lol its on my homework

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

yeaa and my teacher isnt the best

OpenStudy (anonymous):

@evan97 @pretyroze A cubic polynomial has 3 roots, but not necessarily 3 x-intercepts. Example...the roots may be x = 3, 2+3i, and 2 - 3i. You can't plt the two imginary numbers on the real axis, so that polynomial will have one x-intercept.

OpenStudy (anonymous):

In polynomials, you have to be careful not to equate roots with x-intercepts.

OpenStudy (anonymous):

Yes but you can't factor a cubic function given only 1 real root correct?

OpenStudy (anonymous):

but not all cubic polynomials have imaginary numbers correct?

OpenStudy (anonymous):

Yes, you can.

OpenStudy (anonymous):

Correct, the 3 roots may be either all 3 real numbers or 1 real number and 2 imaginary numbers. No other combination is possible for a cubic polynomial.

OpenStudy (anonymous):

so in my case what would it be with an x intercept of 4

OpenStudy (anonymous):

Yeah I'm wondering the same thing

OpenStudy (anonymous):

@pretyroze @evan97 As an exmaple, the equation x^3 - 10x^2 + 37x - 57 = 0 has one real root and two imaginary roots. And this polynomial is factorable into (x-4)(x^2 -6x + 13). The real root is x = 4; the imaginary roots are 3+2i and 3 - 2i.

OpenStudy (anonymous):

But what if an equation is not given. For example, "a cubic function a real zero 4, what is the function"

OpenStudy (anonymous):

That question is impossible correct?

OpenStudy (anonymous):

Which is what his homework is asking right?

OpenStudy (anonymous):

ughhh i give up :(

OpenStudy (anonymous):

i understand the first 2 but this last one is confusing

OpenStudy (anonymous):

Just given the graph with one x-intercept and no equation, there is no way to write the function, unless you know all the roots, or additional clues to the given polynomial.

OpenStudy (anonymous):

ok so i'll just write no solution! :)

OpenStudy (anonymous):

That's what I thought. So than if that's what the question asks, just make up 2 other zeros

OpenStudy (anonymous):

@pretyroze What is confusing? I will try to unconfuse you, so that it is crystal clear.

OpenStudy (anonymous):

So if you have 4, then (x-4)(x-?)(x-?)

OpenStudy (anonymous):

@evan97 Please type your entire question. I dont want to second-guess as to whats bothering you.